OpenStudy (anonymous):

Which is a solution of the equation? 2x - y = 0 (Points : 4) (1, 1) (2,-4) (1, 2) (3, -1)

OpenStudy (anonymous):

Hi

OpenStudy (anonymous):

Can you help me

OpenStudy (anonymous):

Hi

OpenStudy (anonymous):

I need help on this problem

OpenStudy (anonymous):

I am in special ED so be gentle with me O:

OpenStudy (anonymous):

I cant function problems right

OpenStudy (anonymous):

You need to plug in those points for x, y and see which one equal zero

OpenStudy (anonymous):

(2, -4)

OpenStudy (anonymous):

Some how i got this?

OpenStudy (anonymous):

Not quite. (2, -4) would equal 8

OpenStudy (anonymous):

Is it correct

OpenStudy (anonymous):

You have to plug it in right

OpenStudy (anonymous):

I pluged in the 2 into the x

OpenStudy (anonymous):

and -4 into the y

OpenStudy (anonymous):

2*2 is f right

OpenStudy (anonymous):

4-

OpenStudy (anonymous):

And i did 4-4 and i got zero

OpenStudy (anonymous):

Yes you have toplug it in. (2, -4) would become 2(2)-(-4)=4+4=8

OpenStudy (anonymous):

0=0 RIght

OpenStudy (anonymous):

Dont forget that negative * negative equals positive

OpenStudy (anonymous):

Try plugging in points (1, 2)

OpenStudy (anonymous):

No because the there no negative in this problem except the y

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

It equals zero!

OpenStudy (anonymous):

0=0

OpenStudy (anonymous):

So 1,2 is my number set

OpenStudy (anonymous):

(1, 2) is your correct number set. Just try not to get confused next time by the negatives

OpenStudy (anonymous):

Ok